tickets to the local circus cost 3 dollars for children and 5 dollars for adult. There were 3 times as many children tickets sold as adult tickets. All together the circus made 700 dollars. How many children and how many adult bought tickets to the circus?
Answer by Maths68(1474)

Adults tickets sold = x
Children tickets sold= 3x
Total Amount = $700
(Cost of a children ticket)(Children tickets sold)+(Cost of an adult ticket)( Adult tickets sold)= Total Amount
3(3x)+5(x)=700
9x+5x=700
14x=700
14x/14=700/14
x=50


Adults tickets sold = x = 50
Children tickets sold= 3x = 3*50 = 150



Check
=====
3*150+5*50=700
450+250=700
700=700